OpenStudy (anonymous):

The perimeter of a rectangle is 26ft. The length is 7 ft longer than the width. Find the dimensions. Write a system of linear equations and solve the resulting system. Let x be the length and y be the width. 1st equation is 2x+2y=___. 2nd equation is x= y+___. What's the length_____ and width___?

OpenStudy (anonymous):

1st equation : 2x + 2y = 26 (because perimeter is all the sides added together and is 26.) 2nd equation: x=y+7 (because length is 7ft longer than width) Put x form second equation into the 1st eq. 2(y+7) + 2y =26 2y +14 +2y = 26 4y + 14 =26 4y = 12 y = 3 put this into 2nd eq. x = y +7 x = 3 +7 =10

OpenStudy (anonymous):

It tells you that the length 'x' is 7 ft longer than the width 'y'. So you can then form the equation x= y+7